OpenStudy (anonymous):

WIll fan and medal please Given the function f(x) = 0.3(4)x, what is the value of f−1(6)?

OpenStudy (anonymous):

hard to read is it \[f(x)=0.3\times 4^x\]?

OpenStudy (anonymous):

yes and the other side is the same but only -1 is an exponent 6 regular

OpenStudy (anonymous):

lol regular 6 do you know what \(f^{-1}(x)\) means?

OpenStudy (anonymous):

"no" is a fine answer, i am just asking

OpenStudy (anonymous):

no I only know how to do things like f(x)-g(x) this is new

OpenStudy (anonymous):

ok \(f^{-1}\) is the "inverse function" so if for example \(f(2)=7\) then \(f^{-1}(7)=2\)

OpenStudy (anonymous):

more generally if \(f(x)=y\) then \(f^{-1}(y)=x\)

OpenStudy (anonymous):

this one is a kind of hard one to solve, because what you need to do is solve \[0.3\times 4^x=6\]for \(x\)

OpenStudy (anonymous):

so 6-0.3 normally if normally x I'd divide the 4 next

OpenStudy (anonymous):

first step is to divide by \(.3\) and get \[4^x=\frac{6}{.3}=\frac{60}{3}=20\]

OpenStudy (anonymous):

Not minus then

OpenStudy (anonymous):

oh duh its time sorry forgot that

OpenStudy (anonymous):

lets go slow and i want to make sure that it is \[\huge f(x)=0.3\times 4^x\]right? the \(x\) is up in the exponent, correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so we are going to solve \[.3\times 4^x=6\] divide by \(.3\) and get \[4^x=20\] ok with that?

OpenStudy (anonymous):

yeah I understand that part now

OpenStudy (anonymous):

now i am not sure what class you are taking, but you cannot use algebra to solve \[4^x=20\] you have to use logarithms

OpenStudy (anonymous):

are you using logs in your class? perhaps a "pre-calculus" course?

OpenStudy (anonymous):

4x=20 = x = ln (20) over ln (4)

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

= 2.1 right?

OpenStudy (anonymous):

thank you very much! It's much easier explained out now

OpenStudy (anonymous):

looks good http://www.wolframalpha.com/input/?i=ln%2820%29%2Fln%284%29 but if you want to round, use \(2.2\) or \(2.16\)

OpenStudy (anonymous):

yw